What is the Kinetic Energy of a 1300 kg object that is moving with a speed of 30 m/s? *


1
Expert's answer
2022-02-28T13:52:24-0500

The kinetic energy

"E_k=\frac{mv^2}{2}\\
=\frac{1300*30^2}{2}=585000\:\rm J"